OpenStudy (anonymous):

88% of what number is 66? A. 75 B. 58.08 C. 1.3 D. 0.75

OpenStudy (anonymous):

66/88% = 75

OpenStudy (akashdeepdeb):

Let that number be 'NUMBER' So 88% meaning 88 parts in 100 parts of NUMBER = 66 So \[\frac{88 * NUMBER}{100} = 66\] Thus, \[NUMBER = \frac{66 * 100}{88}\] Now calculate! Understood? :)

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

100%

OpenStudy (anonymous):

Convert the decimal to a percent. 3.45 A. 345% B. 34.5% C. 3.45% D. 0.345%

OpenStudy (mathstudent55):

Since this is a multiple choice question, with the choices you are given, you don't even need to calculate it. Just reason it out. A number is 100% of that number. 88% of a number is less than the number since 88% is less than 100%. If 88% of a number is 66, that means the number must be greater than 66. There is only one choice that is greater than 66. The answer must be 75.